# OW-Rep: Open World Object Detection with Instance Representation Learning

### v2

@@ Abstract
Open World Object Detection (OWOD) addresses realistic scenarios where unseen object classes emerge, enabling detectors trained on known classes to detect unknown objects and incrementally incorporate the knowledge they provide. While existing OWOD methods primarily focus on detecting unknown objects, they often overlook the rich semantic relationships between detected objects, which are essential for scene understanding and applications in open-world environments (e.g., open-world tracking and novel class discovery). In this paper, we extend the OWOD framework to jointly detect unknown objects and learn semantically rich instance embeddings, enabling the detector to capture fine-grained semantic relationships between instances. To this end, we propose two modules that leverage the rich and generalizable knowledge of Vision Foundation Models (VFM). First, the Unknown Box Refine Module uses instance masks from the Segment Anything Model to accurately localize unknown objects. The Embedding Transfer Module then distills instance-wise semantic similarities from VFM features to the detector’s embeddings via a relaxed contrastive loss, enabling the detector to learn a semantically meaningful and generalizable instance feature. Extensive experiments show that our method significantly improves both unknown object detection and instance embedding quality, while also enhancing performance in downstream tasks such as open-world tracking.
@@ 1 Introduction
Humans naturally detect and localize unfamiliar objects, drawing on associations with similar ones to comprehend them. While deep learning-based methods have made significant improvements in perception, object detectors still struggle to detect objects beyond the training distribution. To operate reliably in real-world environments, detectors should be able to detect and identify both the previously known and novel unknown objects. Open World Object Detection (OWOD) [18], recently introduced, enhanced the detectors’ practicality and reliability by enabling the detection of unknown objects not labeled in the training set. The framework then involves a human annotator to label detected unknowns and incrementally trains the detector on this new data, enabling continuous adaptation in open-world environments.
@@
Figure 1: We propose a method for training an open-world object detector that not only detects unknown objects but also learns semantically rich feature embeddings that capture meaningful inter-object relationships. Existing OWOD methods [18, 13, 50, 8] primarily focus on detecting unknown objects but overlook the semantic relationships between proposals. Our approach explicitly captures these relationships by enhanced instance embeddings through VFM [34, 22] distillation, while also improving unknown object detection.
@@
To achieve a comprehensive understanding of a scene in open-world environments, it is crucial to not only detect both known and unknown objects but also to learn the finer relationships between them. These relationships play a key role in diverse applications of detectors where predefined information about novel classes is unavailable. For example, tasks such as object tracking [30, 35, 44], path planning [17], and class discovery [46, 11] in open-world scenarios heavily rely on semantic similarity between detected objects for association and obstacle identification. However, existing OWOD approaches primarily focus on learning generalized object representations across diverse categories, often overlooking these finer relationships and failing to capture meaningful semantic information. As a result, methods relying on pretrained detector features to infer relationships between known and unknown objects often struggle to work effectively [37, 14].
@@
Since existing OWOD methods lack supervision to learn rich features of unknown objects, a few studies have tried to enhance the feature quality by self-supervised learning using unknown proposals [46, 11, 47]. However, the inaccurate proposals from detectors hinder the methods of learning robust features. Moreover, these methods mainly focus on learning representations of unknown objects present in the training data, which we refer to as known-unknowns. In real-world deployment, systems may also encounter unknown-unknown objects that were never observed during training [7]. Therefore, the learned feature space should be able to embed both of them properly based on their semantics for robust operation in open-world scenarios.
@@
In this paper, we extend the OWOD framework to simultaneously detect unknown objects and extract semantically rich features, and propose a training method to achieve this (Fig. 1). To enable accurate localization and the learning of semantically rich feature embeddings, our approach leverages the knowledge of Vision Foundation Models (VFMs). Specifically, segmentation masks from Segment Anything Model (SAM) [22] are first utilized to supervise the localization of unknown objects. Moreover, the detector’s instance features are enhanced by distilling the similarity between instances obtained from VFM’s pixel-level features [34] to the detector. This distillation is performed using a relaxed contrastive loss [21], which provides a rich supervisory signal that enables the learning of a generalizable feature space.
@@
The efficacy of the proposed method is demonstrated through extensive experiments, including its application to downstream tasks. Our method significantly enhances unknown detection performance and instance embedding quality. Evaluation with a novel benchmark with unknown-unknowns shows the generalizability of our method in properly embedding both the known-unknown and unknown-unknowns. Furthermore, through an application to open-world tracking, we demonstrate that learning semantically rich instance features benefits downstream tasks.
@@ 2 Related Works
Open World Object Detection. The Open World Object Detection (OWOD) framework is introduced to detect objects in realistic environments where previously unseen objects may appear, allowing the detector to progressively evolve by incorporating newly detected objects [18]. To identify novel objects not labeled in the training set, existing methods learn a generalized notion of object by utilizing the knowledge from labeled, known objects [18, 13, 50, 49, 19, 43, 42, 31, 27, 8, 40]. Some methods directly classify the proposals as either known or unknown using energy functions or activation scores [18, 13], while PROB [50] learns an objectness score based on the distribution of known object features. Hyp-OW [8] further grouped known objects into superclasses and computed objectness based on these groupings. However, as these methods focus on learning a generalized concept of unknown objects, they struggle to distinguish between different unknown objects.
@@
For OWOD to be applicable in diverse real-world scenarios, it must not only detect unknown objects but also capture detailed relationships between them. In particular, open-world tracking [30, 35] associates objects across frames based on semantic similarities between pairwise proposals. Similarly, path planning and mapping often require determining which known class an unknown object mostly resembles to predict its motion or refine maps [39]. Novel class discovery also depends on proposal similarities to identify new categories [46, 11, 14, 37]. Although some methods acquire such relationships by clustering fixed detector features [14, 37], these groupings tend to be coarse, and the features themselves lack meaningful information. To capture finer relationships, the detector must learn semantically rich instance features directly.
@@
Figure 2: Overall Architecture of the Proposed Method. Our method extends OWOD by not only detecting unknown objects but also extracting semantically rich features. We adopt PROB [50] as the baseline open-word object detector. During training, the known and unknown proposals from PROB, with corresponding instance embeddings, are fed into the proposed modules. The Unknown Box Refine Module improves the localization of unknown objects by treating refined unknown boxes from SAM [22] as pseudo ground truth. The Embedding Transfer Module extracts source embeddings by average pooling DINOv2 [34] features within the refined unknown and known proposals. Pairwise similarities between source embeddings are then computed and used as weights for the relaxed contrastive loss [21], controlling the attraction and repulsion between instance embeddings. At inference, the detector generates semantically rich instance embeddings, capturing fine-grained relationships between detected proposals.
@@
Self-Supervised Representation Learning. Due to the lack of supervision for unknown objects, open-world detection methods leveraged self-supervised learning techniques [4, 33, 15, 21, 36] to obtain rich feature embeddings for each instance. OSODD [47] trains a separate proposal embedding network via contrastive learning on augmented pairs of cropped proposals [15]. In contrast, some approaches directly learn instance-level features from the detector itself. UC-OWOD [46] strengthens relationships between highly similar or dissimilar proposals through contrastive learning, while RNCDL [11] clusters all proposals into a fixed number of classes and treats those within the same cluster as positive pairs [1, 6]. However, the effectiveness of these methods depends heavily on the quality of unknown proposals, which is often degraded by inaccurate boxes and high false positive rates in open-world detectors [18, 13, 50]. Furthermore, these methods use binary labels for contrastive loss to enforce unknown class separation, failing to capture finer relationships between instances [21]. To learn richer, more robust features for OWOD, a more informative supervisory signal is required.
@@
Foundation Model for Generalizing Object Detection. Recent efforts have explored using foundation models for unknown object detection due to their strong generalization capabilities in unseen scenarios [32, 49, 5, 2, 41, 20, 25, 29]. However, these models are computationally expensive, as they rely on large-scale image-text or instance-text datasets and often require models during inference. To mitigate this, several approaches focus on distilling knowledge from foundation models into object detectors [10, 26, 16, 12, 23, 45, 24]. For example, instance masks generated by SAM can improve unknown object localization [16]. Some methods also distill features from vision-language models to learn unknown object attributes and improve detection performance [12, 23, 45, 24]. However, these approaches can only detect unknown objects that match predefined text prompts. As a result, they fail to detect truly novel objects outside the prompt set and cannot support incremental learning based on unknown detection results.
@@ 3.2 Overall Architecture
The overall architecture of the proposed method is illustrated in Fig 2. We adopt PROB [50] as the baseline open-world object detector. Built on deformable DETR [48], PROB generates query embeddings that encapsulate information about each predicted object. These embeddings are then processed to compute the bounding boxes and classification scores, where each proposals are classified as either a known object or background. Known proposals are matched with ground truth objects via Hungarian matching, and their regression and classification loss, $\mathcal{L}_{b}$ and $\mathcal{L}_{c}$ [50], are computed. For background proposals, unknown objects are identified based on an objectness score, representing the probability that a proposal corresponds to an object. This score is learned through the objectness loss, $\mathcal{L}_{o}$ [50].
Based on PROB, we propose two modules that are jointly trained with the detector in an end-to-end manner to distill rich VFM features, enabling accurate object detection and semantically rich feature extraction. First, the Unknown Box Refine Module improves the localization of unknown objects by leveraging instance masks generated by SAM [22]. Then, the Embedding Transfer Module distills the instance-wise semantic similarities obtained from the VFM pixel-level features [34] to enhance the detector’s instance embeddings. Both modules are used only during training and discarded at inference.
@@ 3.3 Unknown Box Refine Module
We employ SAM’s high-quality instance masks [22] to supervise the localization of unknown objects and obtain accurate bounding boxes. Given an input image $I$, we select the top-$k$ proposals with the highest objectness socres, excluding those classified as known objects. The bounding boxes of these proposals, $\mathbf{B}_{u}=\{\mathbf{b}_{\text{u},i}\}_{i=1}^{k}$, serve as prompts for SAM to generate refined unknown boxes, $\mathbf{b}_{\text{u},i}^{\text{SAM}}$. To filter out false positives, only refined boxes with an Intersection over Union (IoU) above a threshold $\kappa$ with their corresponding original boxes are retained. The regression head is then trained using these refined boxes as ground truth, and the unknown regression loss calculated as follows:
@@
where $U$ denotes the set of indices corresponding to the retained boxes with sufficiently high overlap, and GIoU represents the Generalized Intersection over Union [38].
Table 1: Quantitative results on OWOD split. We compared our method with the other OWOD methods that learn instance embeddings, reporting performance on Unknown Recall, Recall@1, and known mAP. Our method outperforms other methods in both unknown object detection (Unknown Recall) and feature embedding quality (Recall@1) while maintaining competitive performance on known objects (Known mAP), demonstrating its effectiveness for OWOD and instance representation learning. Note that known mAP is calculated over all known objects. Since all classes are known in Task 4, Unknown Recall and Recall@1 are not reported.
@@
PROB[50] 18.84 4.42 58.98 16.04 3.79 43.76 17.28 6.69 35.81 31.72
OSODD[47] 18.84 4.85 58.98 16.04 3.99 43.76 17.28 6.57 35.81 31.72
UC-OWOD[46] 12.87 3.84 45.11 10.04 2.85 34.04 13.31 4.83 29.30 31.72
RNCDL[11] 18.93 1.41 57.51 15.13 1.38 43.90 17.35 2.82 35.54 31.49
@@ 3.4 Embedding Transfer Module
We utilize the rich feature representations from DINOv2 [34] as the source embedding. To distill this embedding space, a set of bounding boxes $\mathbf{B}=\{\mathbf{b}_{i}\}_{i=1}^{N}$ is utilized, comprising both known proposals obtained via Hungarian matching and refined unknown boxes from Sec. 3.3, where $N$ denotes the total number of such boxes. First, the DINOv2 embedding $\mathbf{F}\in\mathbb{R}^{D}$ is extracted from the input image and resized to its original size. Proposal-wise source embeddings $\mathbf{S}=\{\mathbf{s}_{i}\}_{i=1}^{N}$, each with dimension $D$, are then computed by applying average pooling to the resized DINOv2 feature within each bounding box. Subsequently, the detector’s proposal-wise instance embeddings $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{N}$, each with dimension $d$, are obtained by passing the query embeddings corresponding to $\mathbf{B}$ through an MLP head.
To distill the inter-sample relations of source embeddings to the detector’s embedding, the semantic similarity of pairwise source embeddings is computed from their Euclidean distance, using a Gaussian kernel [21] as follows:
@@
where $\sigma$ is the kernel bandwidth. The instance embeddings are trained using the following relaxed contrastive loss [21]:

### v3

@@ Abstract
Open World Object Detection (OWOD) addresses realistic scenarios where unseen object classes emerge, enabling detectors trained on known classes to detect unknown objects and incrementally incorporate the knowledge they provide. While existing OWOD methods primarily focus on detecting unknown objects, they often overlook the rich semantic relationships between detected objects, which are essential for scene understanding and applications in open-world environments (e.g., open-world tracking and novel class discovery). In this paper, we extend the OWOD framework to jointly detect unknown objects and learn semantically rich instance embeddings, enabling the detector to capture fine-grained semantic relationships between instances. To this end, we propose two modules that leverage the rich and generalizable knowledge of Vision Foundation Models (VFMs) and can be integrated into open-world object detectors. First, the Unknown Box Refine Module uses instance masks from the Segment Anything Model to accurately localize unknown objects. The Embedding Transfer Module then distills instance-wise semantic similarities from VFM features to the detector’s embeddings via a relaxed contrastive loss, enabling the detector to learn a semantically meaningful and generalizable instance feature. Extensive experiments show that our method significantly improves both unknown object detection and instance embedding quality, while also enhancing performance in downstream tasks such as open-world tracking.
@@ 1 Introduction
Humans naturally detect and localize unfamiliar objects, drawing on associations with similar ones to comprehend them. While deep learning-based methods have made significant improvements in perception, object detectors still struggle to detect objects beyond the training distribution. To operate reliably in real-world environments, detectors should be able to detect and identify both the previously known and novel unknown objects. Open World Object Detection (OWOD) [joseph2021towards], recently introduced, enhanced the detectors’ practicality and reliability by enabling the detection of unknown objects not labeled in the training set. The framework then involves a human annotator to label detected unknowns and incrementally trains the detector on this new data, enabling continuous adaptation in open-world environments.
@@
Figure 1: We propose a method for training an open-world object detector that not only detects unknown objects but also learns semantically rich feature embeddings that capture meaningful inter-object relationships. Existing OWOD methods [joseph2021towards, gupta2022ow, zohar2023prob, doan_2024_HypOW] primarily focus on detecting unknown objects but overlook the semantic relationships between proposals. Our approach explicitly captures these relationships by enhanced instance embeddings through VFM [oquab2023dinov2, kirillov2023segment] distillation, while also improving unknown object detection.
@@
To achieve a comprehensive understanding of a scene in open-world environments, it is crucial to not only detect both known and unknown objects but also to learn the finer relationships between them. These relationships play a key role in diverse applications of detectors where predefined information about novel classes is unavailable. For example, tasks such as object tracking [liu2022opening, pang2021quasi, wojke2017simple], path planning [honda2024replan], and class discovery [wu2022uc, fomenko2022learning] in open-world scenarios heavily rely on semantic similarity between detected objects for association and obstacle identification. However, existing OWOD approaches primarily focus on learning generalized object representations across diverse categories, often overlooking these finer relationships and failing to capture meaningful semantic information. As a result, methods relying on pretrained detector features to infer relationships between known and unknown objects often struggle to work effectively [rambhatla2021pursuit, hayes2024pandas].
@@
Since existing OWOD methods lack supervision to learn rich features of unknown objects, a few studies have tried to enhance the feature quality by self-supervised learning using unknown proposals [wu2022uc, fomenko2022learning, zheng2022towards]. However, the inaccurate proposals from detectors hinder the methods of learning robust features. Moreover, these methods mainly focus on learning representations of unknown objects present in the training data, which we refer to as known-unknowns. In real-world deployment, systems may also encounter unknown-unknown objects that were never observed during training [dhamija2020overlooked]. Therefore, the learned feature space should be able to embed both of them properly based on their semantics for robust operation in open-world scenarios.
@@
In this paper, we extend the OWOD framework to simultaneously detect unknown objects and extract semantically rich features (Fig. 1) by introducing modules that can be jointly trained with a baseline open-world object detector. To enable accurate localization and the learning of semantically rich feature embeddings, our approach leverages the knowledge of Vision Foundation Models (VFMs). Specifically, segmentation masks from Segment Anything Model (SAM) [kirillov2023segment] are employed to supervise the localization of unknown objects, while the detector’s instance features are enriched by distilling the similarity relations derived from VFM’s pixel-level features [oquab2023dinov2]. This distillation is performed using a relaxed contrastive loss [kim2021embedding], which provides a rich supervisory signal that enables the learning of a generalizable feature space. Our contribution is the joint integration of box refinement and relaxed contrastive transfer, which allows the detector to learn semantically rich features despite the lack of direct supervision and the ambiguity of class boundaries in open-world scenarios.
@@
We demonstrate the effectiveness of our method through extensive experiments, including downstream applications. Compared to prior self-supervised approaches, our method significantly improves unknown object detection and instance embedding quality. Performance gains across various baseline open-world detectors further validate its strong compatibility. Evaluation with a novel benchmark with unknown-unknowns shows the generalizability of our method in properly embedding both the known-unknown and unknown-unknowns. Finally, through an application to open-world tracking, we demonstrate that learning semantically rich instance features benefits downstream tasks.
@@ 2 Related Works
Open World Object Detection. The Open World Object Detection (OWOD) framework is introduced to detect objects in realistic environments where previously unseen objects may appear, allowing the detector to progressively evolve by incorporating newly detected objects [joseph2021towards]. To identify novel objects not labeled in the training set, existing methods learn a generalized understanding of object by utilizing the knowledge from labeled, known objects [joseph2021towards, gupta2022ow, zohar2023prob, zohar2023open, kim2022learning, wang2023detecting, wang2023random, ma2021annealing, liang2023unknown, doan_2024_HypOW, sun2024exploring]. Some methods directly classify the proposals as either known or unknown using energy functions or activation scores [joseph2021towards, gupta2022ow], while PROB [zohar2023prob] learns an objectness score based on the distribution of known object features. Hyp-OW [doan_2024_HypOW] further grouped known objects into superclasses and computed objectness based on these groupings. However, as these methods focus on learning a generalized concept of objects, they struggle to distinguish between different unknown objects.
@@
For OWOD to be applicable in diverse real-world scenarios, it must not only detect unknown objects but also capture detailed relationships between them. In particular, open-world tracking [liu2022opening, pang2021quasi] associates objects across frames based on semantic similarities between pairwise proposals. Similarly, path planning and mapping often require determining which known class an unknown object mostly resembles to predict its motion or refine maps [sodano2024open]. Novel class discovery also depends on proposal similarities to identify new categories [wu2022uc, fomenko2022learning, hayes2024pandas, rambhatla2021pursuit]. Although some methods acquire such relationships by clustering fixed detector features [hayes2024pandas, rambhatla2021pursuit], these groupings tend to be coarse, and the features themselves lack meaningful information. To effectively capture detailed relationships, the detector must learn semantically rich instance features directly.
@@
Figure 2: Overall Architecture of the Proposed Method. Our method extends OWOD by not only detecting unknown objects but also extracting semantically rich features. We adopt PROB [zohar2023prob] as the baseline detector, but any other open-world object detector can be utilized as a baseline. During training, the known and unknown proposals from PROB, with corresponding instance embeddings, are fed into the proposed modules. The Unknown Box Refine Module improves the localization of unknown objects by treating refined unknown boxes from SAM [kirillov2023segment] as pseudo ground truth. The Embedding Transfer Module extracts source embeddings by average pooling DINOv2 [oquab2023dinov2] features within the refined unknown and known proposals. Pairwise similarities between source embeddings are then computed and used as weights for the relaxed contrastive loss [kim2021embedding], controlling the attraction and repulsion between instance embeddings. At inference, the detector generates semantically rich instance embeddings, capturing fine-grained relationships between detected proposals.
@@
Self-Supervised Representation Learning. Due to the lack of supervision for unknown objects, open-world detection methods have employed self-supervised learning techniques [chen2020simple, oord2018representation, he2020momentum, kim2021embedding, passalis2018learning] to obtain rich feature embeddings for each instance. OSODD [zheng2022towards] trains a separate proposal embedding network through contrastive learning on augmented pairs of cropped proposals [he2020momentum]. In contrast, some approaches directly learn instance-level features from the detector itself. UC-OWOD [wu2022uc] strengthens relationships between highly similar or dissimilar proposals through contrastive learning, while RNCDL [fomenko2022learning] clusters all proposals into a predefined number of classes and treats those within the same cluster as positive pairs [asano2019self, cuturi2013sinkhorn]. However, the effectiveness of these methods depends heavily on the quality of unknown proposals, which is often degraded by inaccurate boxes and high false positive rates in open-world detectors [joseph2021towards, gupta2022ow, zohar2023prob]. Furthermore, these methods use binary labels for contrastive loss to enforce unknown class separation, thus failing to capture detailed relationships between instances [kim2021embedding]. To learn richer and more robust features for OWOD, a more informative and accurate supervisory signal is required.
@@
Foundation Model for Generalizing Object Detection. Recent efforts have explored using foundation models for unknown object detection due to their strong generalization capabilities in previously unseen scenarios [minderer2022simple, zohar2023open, cheng2024yolo, cao2024coda, wang2024ov, kim2024region, li2022grounded, liu2024grounding]. However, these models are computationally expensive, as they rely on large-scale image-text or instance-text datasets and often require models during inference. To mitigate this, several approaches focus on distilling knowledge from foundation models into object detectors [faber2024leveraging, li2024matching, he2023usd, gu2021open, kuo2022f, wu2023aligning, li2023distilling]. For example, instance masks generated by SAM can improve unknown object localization [he2023usd]. Some methods also distill feature from vision-language models to learn unknown object attributes and improve detection performance [gu2021open, kuo2022f, wu2023aligning, li2023distilling, wang2023open, wang2023learning, thawakar2025video, zang2022open]. However, these approaches can only detect unknown objects that match predefined text prompts. As a result, they fail to detect truly novel objects outside the prompt set and cannot support incremental learning based on the detection results of such objects.
@@ 3.2 Overall Architecture
The overall architecture of the proposed method is illustrated in Fig 2. We adopt PROB [zohar2023prob] as the baseline detector and extend it through our introduced modules. Note that any other open-world object detector can also be utilized as a baseline detector. Built on deformable DETR [zhu2020deformable], PROB generates query embeddings that encapsulate information about each predicted object. These embeddings are then processed to compute the bounding boxes and classification scores, where each proposals are classified as either a known object or background. Known proposals are matched with ground truth objects via Hungarian matching, and their regression and classification loss, $\mathcal{L}_{b}$ and $\mathcal{L}_{c}$ [zohar2023prob], are computed. For background proposals, unknown objects are identified based on an objectness score, representing the probability that a proposal corresponds to an object. This score is learned through the objectness loss, $\mathcal{L}_{o}$ [zohar2023prob].
Based on PROB, we propose two modules that are jointly trained with the detector in an end-to-end manner to distill rich VFM features, enabling accurate object detection and semantically rich feature extraction. First, the Unknown Box Refine Module improves the localization of unknown objects by leveraging instance masks generated by SAM [kirillov2023segment]. Then, the Embedding Transfer Module distills the instance-wise semantic similarities obtained from the VFM pixel-level features [oquab2023dinov2] to enhance the detector’s instance embeddings. Both modules are used only during training and discarded at inference.
@@ 3.3 Unknown Box Refine Module
We employ SAM’s high-quality instance masks [kirillov2023segment] to supervise the localization of unknown objects and obtain accurate bounding boxes. Given an input image $I$, we select the top-$k$ proposals with the highest objectness scores, excluding those classified as known objects. The bounding boxes of these proposals, $\mathbf{B}_{u}=\{\mathbf{b}_{\text{u},i}\}_{i=1}^{k}$, serve as prompts for SAM to generate refined unknown boxes, $\mathbf{b}_{\text{u},i}^{\text{SAM}}$. To filter out false positives, only refined boxes with an Intersection over Union (IoU) above a threshold $\kappa$ with their corresponding original boxes are retained. The regression head is then trained using these refined boxes as ground truth, and the unknown regression loss calculated as follows:
@@
where $U$ denotes the set of indices corresponding to the retained boxes with sufficiently high overlap, and GIoU represents the Generalized Intersection over Union [rezatofighi2019generalized].
Table 1: Quantitative results on OWOD split. We compared our method with the other approaches that simultaneously detect unknown objects and learn instance embeddings, reporting performance on Unknown Recall, Recall@1, and known mAP. Our method outperforms others in both unknown object detection (Unknown Recall) and feature embedding quality (Recall@1) while maintaining competitive performance on known objects (Known mAP), demonstrating its effectiveness for OWOD and instance representation learning. Note that known mAP is calculated over all known objects. Since all classes are known in Task 4, Unknown Recall and Recall@1 are not reported.
@@
PROB[zohar2023prob] 18.84 4.42 58.98 16.04 3.79 43.76 17.28 6.69 35.81 31.72
OSODD[zheng2022towards] 18.84 4.85 58.98 16.04 3.99 43.76 17.28 6.57 35.81 31.72
UC-OWOD[wu2022uc] 12.87 3.84 45.11 10.04 2.85 34.04 13.31 4.83 29.30 31.72
RNCDL[fomenko2022learning] 18.93 1.41 57.51 15.13 1.38 43.90 17.35 2.82 35.54 31.49
@@ 3.4 Embedding Transfer Module
We utilize the rich feature representations from DINOv2 [oquab2023dinov2] as the source embedding. To distill this embedding space, a set of bounding boxes $\mathbf{B}=\{\mathbf{b}_{i}\}_{i=1}^{N}$ is utilized, comprising both known proposals obtained via Hungarian matching and refined unknown boxes from Sec. 3.3, where $N$ denotes the total number of such boxes. First, the DINOv2 embedding $\mathbf{F}\in\mathbb{R}^{D}$ is extracted from the input image and resized to its original size. Proposal-wise source embeddings $\mathbf{S}=\{\mathbf{s}_{i}\}_{i=1}^{N}$, each with dimension $D$, are then computed by applying average pooling to the resized DINOv2 feature within each bounding box. Subsequently, the detector’s proposal-wise instance embeddings $\mathbf{Z}=\{\mathbf{z}_{i}\}_{i=1}^{N}$, each with dimension $d$, are obtained by passing the query embeddings corresponding to $\mathbf{B}$ through an MLP head.
To distill the inter-sample relations of source embeddings to the detector’s embedding, the semantic similarity of pairwise source embeddings is computed from their Euclidean distance, using a Gaussian kernel [kim2021embedding] as follows:
@@
where $\sigma$ is the kernel bandwidth. The instance embeddings are trained using the following relaxed contrastive loss [kim2021embedding]: